OpenStudy (anonymous):

One polygon has a side of length 2 feet. A similar polygon has a corresponding side of length 4 feet. The ratio of the perimeter of the smaller polygon to the larger is

OpenStudy (anonymous):

the ratio of the perimeters is the same as the ratio of the sides. smaller/larger = 2/4 = 1/2